### v1

# Fidelity Isn’t Accuracy: When Linearly Decodable Functions Fail to Match the Ground Truth

Abstract
Neural networks excel as function approximators, but their complexity often obscures the nature of the functions they learn. In this work, we propose the linearity score $\lambda(f)$, a simple and interpretable diagnostic that quantifies how well a regression network’s output can be mimicked by a linear model. Defined as the $R^{2}$ between the network’s predictions and those of a trained linear surrogate, $\lambda(f)$ offers insight into the linear decodability of the learned function. We evaluate this framework on both synthetic ($y=x\cdot\sin(x)+\epsilon$) and real-world datasets (Medical Insurance, Concrete, California Housing), using dataset-specific networks and surrogates. Our findings show that while high $\lambda(f)$ scores indicate strong linear alignment, they do not necessarily imply predictive accuracy with respect to the ground truth. This underscores both the promise and the limitations of using linear surrogates to understand nonlinear model behavior—particularly in high-stakes regression tasks.
1 Introduction
Neural networks have revolutionized supervised learning, but their internal complexity often renders their learned functions opaque. In effect, they behave as black boxes. This lack of interpretability hinders trust, accountability, and transparency—especially in high-stakes domains such as healthcare, finance, and criminal justice, where understanding a model’s decisions can be as important as their accuracy. While many techniques exist for interpreting classification networks—such as saliency maps, feature attribution methods like LIME [10] and SHAP [6], and probing approaches [1]—the interpretability of regression networks remains comparatively underexplored, especially at the level of input-output function behavior [6, 14].
One common strategy in model interpretability is to approximate a complex system with a simpler, interpretable surrogate, such as a linear model or decision tree [3, 12]. In classification settings, surrogate models are often applied to internal representations to evaluate the linear accessibility of information at each layer [1, 5]. However, in regression tasks, there has been limited work focused on assessing the linearity of a network’s full input–output function.
In this paper, we introduce a simple yet powerful idea: we quantify how linearly decodable a trained regression network’s output function is. We define a metric, $\lambda(f)$, as the coefficient of determination $R^{2}$ between a neural network’s predictions and those of a linear model trained to mimic it. We refer to this as a measure of how “linearly decodable” the network’s function is—that is, how well its input-output behavior can be approximated by a linear model. This metric offers a scalar summary of how faithfully a linear surrogate can approximate the network’s behavior.
To evaluate the utility and limitations of this approach, we perform experiments on synthetic and real-world regression datasets. For each, we compare a baseline linear model, a trained neural network, and a linear surrogate trained to mimic the network. Our findings show that $\lambda(f)$ provides a reliable measure of how linearly decodable a network’s function is—but also reveal that high fidelity to the network does not always translate into predictive success on the original task.
2 Related Works
A common approach to interpreting neural networks involves training simpler surrogate models—such as linear regressors or decision trees—to approximate their behavior. Notably, Craven and Shavlik [3] introduced TREPAN, an algorithm that uses decision trees to extract disjunctive normal form (DNF) rules by approximating the overall function computed by a neural network. Sato and Tsukimoto [12] proposed CRED, which also employs decision trees but adopts a structural approach: it trains one decision tree to approximate the output layer and additional trees for each neuron in the single hidden layer. While TREPAN operates purely on input-output behavior, CRED utilizes intermediate activations to produce layer-wise symbolic rules. In contrast to both of these algorithms, our method uses a linear surrogate not to extract rules, but to quantify the degree to which a neural network’s output function is linearly decodable.
Another strategy for interpreting neural networks is probing. In this approach, smaller linear models—called “probes”—are trained to predict task-relevant labels from internal representations, providing insight into a model’s training progress or the structure of its learned features. For example, Alain and Bengio [1] introduced linear classifier probes that are inserted into each layer of a trained (or partially trained) neural network to assess how much task-relevant information is linearly accessible at that layer. This method allows researchers to analyze the flow of information during training and characterize the representational capacity of different layers. Hewitt and Manning [5] propose a structural probe to test whether syntactic information is linearly encoded in contextual word embeddings. They learn a linear transformation such that squared distances between transformed word vectors approximate syntactic distances in a dependency parse tree. This approach evaluates whether hierarchical structure is recoverable through linear geometry. While probing methods aim to characterize internal representations, our method treats the network as a black box and applies linear approximation to its overall function rather than examining representations at individual layers.
Most probing methods train predictive models on frozen representations, but Saphra and Lopez [11] take a different approach. They use Singular Vector Canonical Correlation Analysis (SVCCA) [9] to track how linguistic structure emerges over the course of language model training, comparing internal representations to those learned by supervised taggers. Though not a probe in the strict sense, their work shares our goal of tracing when and how interpretable structure emerges. Our method differs by focusing on a concrete, output-level metric of linearity
Other work has analyzed the complexity of neural network functions using measures such as sensitivity, Lipschitz continuity, and generalization bounds [4, 7], offering theoretical insight into expressiveness and stability. Our approach is complementary: rather than analyzing complexity directly or probing internal representations, we treat the network as a black box and ask how much of its behavior is recoverable through linear approximation. By focusing on the output function itself—rather than intermediate layers—we provide a simple, scalable diagnostic that captures one interpretable dimension of the network’s learned behavior.
3 Methodology
To assess whether a trained neural network’s learned function is linearly structured, we propose a projection-based framework that measures how closely its output can be approximated by a linear model. Our goal is not to probe internal layers or compute gradient attributions, but to treat the network as a black-box function $f\colon\mathbb{R}^{n}\to\mathbb{R}$, and ask: to what extent is the behavior of $f$ linearly decodable from its input?
We formalize this question by projecting $f$ onto the space of linear functions and measuring the quality of that approximation.
Let $\mathcal{D}\subset\mathbb{R}^{n}$ be a domain over which the network operates, and assume $f\in L^{2}(\mathcal{D})$, the space of square-integrable functions. Let $\mathcal{L}\subset L^{2}(\mathcal{D})$ denote the subspace of affine functions. We define the optimal linear surrogate as:
$g^{*}=\arg\min_{g\in\mathcal{L}}\mathbb{E}_{x\sim\mathcal{D}}\left[(f(x)-g(x))^{2}\right]$
This is the orthogonal projection of $f$ onto $\mathcal{L}$, and represents the best linear approximation to the network’s function in mean squared error.
We then define the linearity score $\lambda(f)$ as the proportion of variance in the network’s output that is captured by this surrogate:
$\lambda(f):=R^{2}(f,g^{*})=1-\frac{\mathbb{E}[(f(x)-g(x))^{2}]}{\operatorname{Var}(f(x))}$
This score takes values in $(-\infty,1]$, though it typically lies between 0 and 1. When $\lambda(f)\approx 1$, the network’s output function lies close to a linear subspace; when $\lambda(f)\approx 0$, it deviates substantially.
It is important to clarify that $\lambda(f)$ does not measure the linearity of the original data $(x,y)$, but rather the linearity of the network’s learned function $f$ with respect to its input $x$. That is, even if the underlying data is highly nonlinear, a network may learn a function that is approximately linear—making its outputs linearly decodable. Conversely, a network may learn a highly nonlinear function even on linearly structured data.
While the surrogate model $g$ provides a closed-form approximation of $f$, it is not guaranteed to preserve the predictive accuracy of the original network. As our experiments show, even high-fidelity approximations (high $\lambda(f)$) often fail to match the neural network’s performance on the original target. This distinction is central to our analysis.
Ultimately, the primary contribution of this framework is not the surrogate model itself, but the linearity score $\lambda(f)$—a simple, interpretable diagnostic of how linearly structured a network’s output function is.
4 Experiments
To evaluate the proposed framework, we apply it to synthetic and real-world datasets, including the Medical Insurance Cost dataset [2], the Concrete Compressive Strength dataset [13], and the California Housing dataset [8].
For each dataset, we evaluate three models [^1] Experimental code is available on GitHub: https://github.com/jacksoneshbaugh/lambda-linearity-score: (1) a baseline linear regression model trained directly on the input–output pairs; (2) a neural network tailored to the dataset (e.g., multi-layer perceptrons with varying layers, neurons, and ReLU activations, trained with Adam for 50–200 epochs—see Table 1); and (3) a linear surrogate model trained to approximate the output of the neural network. This setup allows us to compute and compare the linearity score $\lambda(f)$ across diverse problem domains, and to assess how closely the neural network’s behavior can be captured by a linear approximation.
Dataset Network Architecture
Synthetic $\mathbf{X}\rightarrow$ 64 $\rightarrow$ ReLU $\rightarrow$ 64 $\rightarrow$ ReLU $\rightarrow$ 1
Medical Costs $\mathbf{X}\rightarrow$ 32 $\rightarrow$ ReLU $\rightarrow$ 16 $\rightarrow$ ReLU $\rightarrow$ 1
Concrete $\mathbf{X}\rightarrow$ 128 $\rightarrow$ ReLU $\rightarrow$ Dropout(0.1) $\rightarrow$ 64 $\rightarrow$ ReLU $\rightarrow$ Dropout(0.1) $\rightarrow$ 32 $\rightarrow$ ReLU $\rightarrow$ 1
California Housing $\mathbf{X}\rightarrow$ 64 $\rightarrow$ ReLU $\rightarrow$ 64 $\rightarrow$ ReLU $\rightarrow$ 1
Table 1: Architectures of the neural networks used in each experiment. Each sequence denotes the number of units, activation functions, and any regularization layers (e.g., Dropout). These networks were selected to balance expressivity and generalization for the respective datasets.
4.1 Synthetic
We begin with a synthetic regression task defined by the function
$y=x\cdot\sin(x)+\epsilon,\text{ where }\epsilon\sim\mathcal{N}(0,0.2^{2})\text{ and }x\in[-4,4]$
This function exhibits a continuous, nonlinear structure, providing a testbed for analyzing the expressiveness of linear approximations to neural models. We additionally introduce Gaussian noise to simulate real-world observational error. Figure 1 illustrates a sample of the dataset.
Figure 1: Synthetic dataset $y=x\cdot\sin(x)+\epsilon$ with Gaussian noise $\epsilon\sim\mathcal{N}(0,0.2^{2})$. The dataset demonstrates a smooth nonlinear relationship.
4.2 Medical Insurance Cost
The Medical Insurance Cost dataset contains diverse profiles of individuals, including features such as age, gender, BMI, number of children, and smoking status. For each profile, the dataset records an associated insurance cost. The relationships among features are moderately nonlinear, particularly for smoking status and age, making this a useful benchmark for testing linear decodability.
4.3 Concrete Compressive Strength
The Concrete Compressive Strength dataset models a nonlinear function from eight features—including cement, water, blast furnace slag, and age of the concrete sample—to compressive strength. Each input exhibits a different degree of nonlinearity with respect to the target, making this dataset well-suited for evaluating the limitations of linear surrogates.
4.4 California Housing
The California Housing dataset includes demographic and geographic information for housing blocks from the 1990 U.S. Census. Features include median income, median house age, average number of rooms and bedrooms, population, average household size, and geographic coordinates. The target variable is median house value (in hundreds of thousands of dollars). With its moderate nonlinearity and blend of structured and unstructured features, this dataset serves as a challenging and realistic case for testing how well the function can be linearly approximated.
5 Results and Discussion
5.1 Results
Across all experiments, we found that the linear surrogate $g$ fails to match the neural network’s predictive accuracy on the original target $y$, even when it closely approximates the network’s output $f$. In every dataset, the surrogate underperforms the neural network in predicting $y$, despite achieving high $\lambda(f)$ in some cases. This indicates that the portions of the function that escape linear approximation—though smaller in variance—are critical for prediction.
In the Medical dataset, for instance, $\lambda(f)\approx 0.93$, yet the surrogate performs worse than the baseline linear model. This suggests that the network’s predictive gains are concentrated in its nonlinear behavior, which $g$ cannot capture. In the Concrete and California datasets, we observe similar effects: the surrogate tracks the neural network well but fails to preserve its advantage. Only in the synthetic dataset is this mismatch more extreme, with low $\lambda(f)$ and complete surrogate failure. Together, these results highlight a key limitation: linear decodability does not guarantee predictive usefulness. Below, we discuss findings—summarized in Tables 2 and 3—on each dataset in detail  [^2] Due to the stochastic nature of gradient descent and nondeterministic hardware behavior, reported values (e.g., $\lambda(f)$, $R^{2}$) may vary slightly across runs. All results shown are representative of the general trends observed across repeated trials..
Dataset Linear R2 NN R2 Mimic R2 $\lambda(f)$
Synthetic -0.0080 0.9755 -0.0137 -0.0105
Medical Costs 0.7836 0.8673 0.7835 0.9186
Concrete 0.5944 0.8588 0.5821 0.6659
California Housing 0.5758 0.7908 0.5658 0.6968
Table 2: Performance metrics across all datasets. The surrogate model’s R2 against the neural network’s output defines the linearity score $\lambda(f)$, reflecting the decodability of the network function.
Dataset NN R2 Mimic R2 $\lambda(f)$
Synthetic 0.9753 $\pm$ 0.0002 -0.0140 $\pm$ 0.0005 -0.0107 $\pm$ 0.0004
Medical Costs 0.8682 $\pm$ 0.0033 0.7832 $\pm$ 0.0003 0.9219 $\pm$ 0.0056
Concrete 0.8610 $\pm$ 0.0073 0.5842 $\pm$ 0.0028 0.6573 $\pm$ 0.0152
California Housing 0.7902 $\pm$ 0.0036 0.5634 $\pm$ 0.0035 0.6952 $\pm$ 0.0017
Table 3: Summary statistics over 5 runs for neural network performance ($R^{2}$), mimic accuracy ($R^{2}$), and linearity score $\lambda(f)$, reported as mean $\pm$ standard deviation.
5.1.1 Synthetic
In the case of the synthetic dataset, our baseline linear model fails to predict the data well, yielding $R^{2}=-0.0080$. The mimic model $g$ performs nearly identically to the baseline linear regression ($R^{2}=-0.0137$), offering no meaningful improvement despite being trained to match the network’s output. Notably, both $R^{2}$ values are less than zero, meaning they both have a worse fit than a constant predictor equal to the mean of the dataset. This failure is expected: the target function is globally nonlinear, with no regions of local linearity for a simple model to exploit.
The neural network itself performs very well, achieving an $R^{2}$ score of $0.9755$. However, the surrogate’s approximation is extremely poor, yielding $\lambda(f)=-0.0105$. This indicates that the learned function $f$ is fundamentally non-linear in input space. That is, no linear transformation $g$ is able to approximate $f$ over the data distribution—even the best available linear fit performs worse than the mean of $f$. Figure 2 visualizes the outputs of all three models.
Figure 2: Predictions of the baseline linear model, neural network, and mimic model on the synthetic dataset.
5.1.2 Medical Insurance Cost
The Medical Insurance Cost dataset presents a moderately structured real-world task. A baseline linear model performs reasonably well $(R^{2}=0.7836)$, while the neural network achieves a higher $R^{2}=0.8673$, indicating its ability to capture additional nonlinear patterns.
Yet despite the neural network’s superior performance, the linear surrogate fails to retain that advantage: it achieves a lower $R^{2}=0.6659$ against the true target, even while attaining a high linearity score $\lambda(f)=0.9186$. This discrepancy reveals a key insight: the linearly decodable portion of the network’s behavior is not necessarily what makes it accurate. The surrogate mimics the network closely, but what it captures may not be what matters.
Figure 3: Predictions from the baseline linear model, neural network, and linear surrogate on the Medical Insurance Cost dataset. Despite closely matching the network $(\lambda(f)=0.9186)$, the surrogate underperforms on the true target.
In short, these results illustrate a core limitation of surrogate-based interpretability: a mimic can generalize to the model without generalizing to the data. Linearity in the learned function does not imply alignment with the data-generating process.
5.1.3 Concrete Compressive Strength
In this dataset, the baseline linear regression model achieves an $R^{2}$ score of 0.5944, reflecting a moderate ability to capture the structure in the data. The neural network substantially improves upon this with $R^{2}=0.8588$, indicating that it successfully models complex nonlinear relationships among the input features.
Figure 4: Predictions from the baseline linear model, neural network, and linear surrogate on the Concrete Compressive Strength dataset. The surrogate achieves moderate fidelity to the neural network ($\lambda(f)=0.6659$), but fails to match its predictive performance.
Figure 4 visualizes the predictions from all three models. The linear surrogate trained to mimic the neural network achieves an $R^{2}$ of 0.5821 against the true target—close to the baseline’s 0.5944—but significantly lower than the network’s 0.8588. Nonetheless, the surrogate yields a linearity score of $\lambda(f)=0.6659$, indicating that it captures over 60% of the variance in the network’s output.
This case again illustrates the gap between linear decodability and predictive utility: although the surrogate mimics much of the network’s behavior, the nonlinear residuals it misses are critical to the network’s superior performance.
5.1.4 California Housing
On this dataset, the baseline linear regression model achieves an $R^{2}$ score of 0.5758, indicating moderate linear structure in the input features. The neural network improves on this with $R^{2}=0.7908$, demonstrating that it captures additional nonlinear patterns in the data. The linear surrogate trained to mimic the neural network achieves $R^{2}=0.5658$ against the true target—slightly worse than the baseline—but exhibits relatively high fidelity to the network, with a linearity score of $\lambda(f)=0.6968$.
Figure 5: Predictions on the California Housing dataset. The linear surrogate tracks the neural network’s output well ($\lambda(f)=0.6968$) but does not preserve its predictive advantage over the baseline.
Figure 5 visualizes predictions from all three models. The linear surrogate closely approximates the neural network’s output but falls short of matching its predictive accuracy on the true target. This gap suggests that while the learned function is largely linearly decodable, its remaining nonlinear components play an outsized role in predictive performance.
As in the Concrete dataset, this result reinforces our central claim: a high $\lambda(f)$ reflects how well a neural network’s behavior can be mimicked, but it does not guarantee that the mimic will retain predictive strength. The predictive signal may be concentrated in the nonlinear residuals that a linear surrogate cannot express.
5.2 Discussion
Our experiments demonstrate that the linearity score $\lambda(f)$ serves as a reliable measure of how linearly decodable a neural network’s output function is. In settings where the learned function is highly nonlinear, such as our synthetic dataset, $\lambda(f)$ drops to zero, confirming that no meaningful linear approximation exists. In contrast, when $\lambda(f)$ is moderate or high—as in the Concrete and California Housing datasets—the surrogate tracks the network’s output with reasonable fidelity, but fails to retain its predictive advantage.
Importantly, a high linearity score indicates fidelity to the neural network, but not necessarily to the ground truth. For example, in the case of the Concrete dataset, approximately 36% of the network’s output variance is not captured by the linear surrogate. However, that residual variance appears to contain nearly all of the model’s predictive improvement over the baseline. The results for the California Housing dataset share this property. These findings suggest that the portion of the neural network’s function that escapes linear approximation—though smaller in variance—may disproportionally contribute to the model’s predictive accuracy.
This distinction highlights an important nuance: a linear surrogate with high $\lambda(f)$ offers insight into the network’s behavior, but may not reproduce its predictive power if that power depends on nonlinear components. In some high-$\lambda$ regimes, the surrogate may still track the network’s output well, but its usefulness as a predictive substitute depends on how much of the model’s performance derives from linearly decodable structure.
While $\lambda(f)$ serves as a useful diagnostic of linear decodability, it is not a comprehensive interpretability method. It offers no insight into a network’s internal mechanisms—such as feature salience, representational geometry, or pathway analysis—nor does it produce explanations in terms of input features, like importance scores or causal attributions. Furthermore, our approach is currently limited to regression tasks; extending it to classification would require adapting the framework to logit or softmax outputs. These limitations suggest that $\lambda(f)$ should be viewed not as a general-purpose interpretability tool, but as a focused, principled diagnostic—particularly well-suited for assessing whether a network’s behavior exhibits linear structure at the output level.
6 Conclusion
Neural networks are often—and justifiably—described as black boxes, posing challenges in high-stakes domains like healthcare and finance. In this work, we proposed a way to measure just how illuminated or comprehensible a network is by quantifying its linear decodability. The metric, $\lambda(f)$, offers a simple and interpretable means of assessing whether a neural network’s output function is well-approximated by a linear model. It is straightforward to compute and opens the door to further analysis of learned functions in terms of their structural simplicity.
Several promising directions emerge from this work. A natural extension would be to explore how an analogous metric might be applied to classification tasks—such as by projecting logits or softmax outputs—to assess interpretability in non-regression settings. Unlike continuous-valued regression outputs, classification involves discrete decision boundaries, which complicates the notion of linear decodability. However, structure in logit space may still reveal interpretable trends—such as class separability or margin alignment—that a linear surrogate could capture.
Another compelling direction is to move beyond quantifying linear decodability and begin characterizing it. Since $\lambda(f)$ captures the proportion of output variance explained by a linear model, a natural next step is to ask: what components of the network’s behavior are responsible for this? Investigating which input regions, features, or latent representations contribute most to the linearly decodable structure could deepen our understanding of how and where simple functional patterns emerge within otherwise complex models. A similar investigation into the components of the network’s behavior that are not linearly representable is equally—if not more—promising.
Across both synthetic and real-world datasets, our experiments show that $\lambda(f)$ reliably captures how linearly decodable a network’s output function is. However—and this is the central finding of this work—high fidelity to the network’s output does not imply high accuracy on the original prediction task. In several cases, linear surrogates with high $\lambda(f)$ scores closely mimic the neural network’s outputs, yet underperform even a baseline linear model when evaluated against the ground truth. This mismatch reveals that the portions of a network’s behavior that are not linearly approximable—though smaller in variance—often carry the critical predictive signal. As a result, $\lambda(f)$ should be interpreted as a measure of structural alignment with the model, not necessarily with the data-generating process.
By providing a concrete measure of when neural network behavior can be meaningfully approximated by simpler linear models, $\lambda(f)$ represents a step toward more interpretable machine learning systems—systems that acknowledge both the power and the opacity of deep networks while offering a principled diagnostic for understanding their behavior.
Acknowledgments
I thank Professor Jorge Silveyra for a valuable discussion that helped spark this project. I am also grateful for the support of the Lafayette College Computer Science Department.